# On Root Capacity, Intersection Indicium, Minimal Generating Sets of Galois Closure & Compositum Feasible Triplets

### v2

@@ 8. Compositum Feasible Triplets
  • compositum feasible over $K$ if there exist $L/K$ an $L^{\prime}/K$ of degrees $a$ and $b$ respectively with the compositum $LL^{\prime}/K$ having degree $c$.
@@
Suppose $c=lcm(a,b)\cdot t$ where t|gcd(a,b)t\ |\ gcd(a,b). Thus $c=ab/s$ where s|gcd(a,b)s\ |\ gcd(a,b). We have $v_{p}(c)=v_{p}(a)+v_{p}(b/s)=v_{p}(a/s)+v_{p}(b)$ for every prime. Hence the triplet clearly satisfies the given condition.
Conversely suppose that the triplet satisfies the condition. Now $a=\Pi p^{v_{p}(a)},b=\Pi p^{v_{p}(b)},$
$c=\Pi p^{v_{p}(c)},lcm(a,b)=\Pi p^{max\{v_{p}(a),v_{p}(b)\}},gcd(a,b)=\Pi p^{min\{v_{p}(a),v_{p}(b)\}}$ and $ab=\Pi p^{v_{p}(a)+v_{p}(b)}$. Since $max\{v_{p}(a),v_{p}(b)\}\leq v_{p}(c)$ for every prime $p$. Hence $lcm(a,b)|c$.
Now $t=c/lcm(a,b)=\Pi p^{v_{p}(c)-max\{v_{p}(a),v_{p}(b)\}}$. Since $v_{p}(c)\leq v_{p}(a)+v_{p}(b)=max\{v_{p}(a),v_{p}(b)\}+min\{v_{p}(a),v_{p}(b)\}$ for every prime number $p$. Thus t|gcd(a,b)t\ |\ gcd(a,b).∎
@@
9. Improving on the Inverse Problems
Lemma 9.1.
Proof.
The following is an improvement on Theorem 5.1.
Theorem 9.2.
Proof.
Remark 9.2.1.
Remark 9.2.2.
Similarly we have an improvement on Theorem 6.1.
Theorem 9.3.
An improvement on Inverse root capacity problem for number fields Theorem 5.2.
Theorem 9.4.
An improvement on Theorem 7.6.
Theorem 9.5.
Let $K$ be a number field and $n>2$ an odd composite number. Then we get arbitrarily large finite families of degree $n$ extensions $L/K$ inside $\bar{K}$ which are pairwise non-isomorphic over $K$ and are pairwise linearly disjoint over $K$, for which the Galois closure has minimal generating sets of cardinalities $2,3,\dots,\omega(n)$ and these are the only possible cardinalities for minimal generating set of Galois closure of $L/K$.
An improvement on Theorem 7.7.
Theorem 9.6.
An improvement on Theorem 8.7.
Theorem 9.7.
Acknowledgements: The author would like to thank Prof Purusottam Rath, CMI Chennai for suggesting to consider the question of isomorphism classes of extensions with given degree and cluster size which is dealt with in Section 9. The author is grateful to Prof P Vanchinathan and Dr Anand Chitrao for suggesting to generalize the main result in [10] for number fields which is achieved in a result in Section 7. The author would also like to thank Dr Shripad Garge, IIT Bombay for introducing the paper by Drungilas et al. [6] and suggesting to investigate the problem of compositum feasible triplets and associated Galois groups which is dealt with in Section 8. Finally the author would like to acknowledge support of IIT Bombay Institute Post Doctoral Fellowship during the later part of this work.

### v3

@@ 8. Compositum Feasible Triplets
  • compositum feasible over $K$ if there exist $L/K$ and $L^{\prime}/K$ of degrees $a$ and $b$ respectively with the compositum $LL^{\prime}/K$ having degree $c$.
@@
Suppose $c=lcm(a,b)\cdot t$ where t|gcd(a,b)t\ |\ gcd(a,b). Thus $c=ab/s$ where s|gcd(a,b)s\ |\ gcd(a,b). We have $v_{p}(c)=v_{p}(a)+v_{p}(b/s)=v_{p}(a/s)+v_{p}(b)$ for every prime. Hence the triplet clearly satisfies the given condition. Conversely suppose that the triplet satisfies the condition. Now $a=\Pi p^{v_{p}(a)},b=\Pi p^{v_{p}(b)},$ $c=\Pi p^{v_{p}(c)},lcm(a,b)=\Pi p^{max\{v_{p}(a),v_{p}(b)\}},gcd(a,b)=\Pi p^{min\{v_{p}(a),v_{p}(b)\}}$ and $ab=\Pi p^{v_{p}(a)+v_{p}(b)}$. Since $max\{v_{p}(a),v_{p}(b)\}\leq v_{p}(c)$ for every prime $p$. Hence $lcm(a,b)|c$. Now $t=c/lcm(a,b)=\Pi p^{v_{p}(c)-max\{v_{p}(a),v_{p}(b)\}}$. Since $v_{p}(c)\leq v_{p}(a)+v_{p}(b)=max\{v_{p}(a),v_{p}(b)\}+min\{v_{p}(a),v_{p}(b)\}$ for every prime number $p$. Thus t|gcd(a,b)t\ |\ gcd(a,b).∎
@@
An improvement on Theorem 8.7 (Similar improvement as in Theorem 5.4).
Theorem 8.8.
@@
Acknowledgements: The author would like to thank Prof Purusottam Rath, CMI Chennai for suggesting to consider the question of isomorphism classes of extensions with given degree and cluster size which is dealt with in Section 5. The author is grateful to Prof P Vanchinathan and Dr Anand Chitrao for suggesting to generalize one of the results in [10] for number fields which is achieved in a result in Section 7. The author would also like to thank Dr Shripad Garge, IIT Bombay for introducing the paper by Drungilas et al. [6] and suggesting to investigate the problem of compositum feasible triplets and associated Galois groups which is dealt with in Section 8. Finally the author would like to acknowledge support of IIT Bombay Institute Post Doctoral Fellowship during the later part of this work.
Data Availability Statement: No new data were created or analyzed in this study.